OpenStudy (theedwardsfamily):

What ratio is equivalent to 14/28? @Legends @Holly00d1248

OpenStudy (theedwardsfamily):

i meant 14/28

OpenStudy (texaschic101):

14/28 = 2/4 = 1/2

OpenStudy (legends):